If the measure of an angle is 78 less than the measure of its compliment, what is the measure of the angle?
Would it be x-78=90?

Answer by drj(1380)

If the measure of an angle is 78 less than the measure of its complement, what is the measure of the angle?

Would it be x-78=90?

Not quite but you are on the right track...here's one approach.

Step 1. Complementary angles means two angles add up to 90 degrees.

Step 2. Let x be the angle.

Step 3. Let 90-x be its complement.

Step 4 x=90-x-78 since the measure of the angle x is 78 less than its complement (90-x).

Step 5. Solving equation in Step 4 yields x=12-x or 2x=12. Then x=6 after dividing by 2. Its complement is 90-x=84 degrees which is 78 degrees less than the complement of 84.

Step 6. ANSWER: The angle is 6 degrees and its complement is 84 degrees.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J